27. The half life for a first order reaction is 5 x 104 s. What percentage of the initial reactant
will react in 2 hours? Calculate
1
Expert's answer
2012-12-19T10:46:19-0500
t(1/2) = 5 x 10^4 - half life for a first orderreaction

t(1/2) = ln2/K
K = ln2/t(1/2)
K = ln2/ 5 x 10^4 = 1.38 x 10^-5

ln(Cx/Co) = -Kt

t = 7200 s

ln(Cx/Co) = -(1.38 x 10^-5 * 7200)

Cx/Co = 0.905 if Co = 1, x = 0.905 (9.05 %)